zarkam21:

Describe the amplitude and frequency of the sound wave for a high, loud note. Identify the characteristic of the sound that is connected to each wave property.

Vocaloid:

amplitude is related to volume frequency is related to pitch for a high + loud note would we expect a high or low amplitude? what about high or low frequency?

zarkam21:

high amplitude and low frequency i think

Vocaloid:

 high amplitude is good, for a high-pitched sound we would expect a high frequency

zarkam21:

For a high loud note wave, there is a high freiquency as well as a high amplitude. The characteristics that are related to these characteristics are; amplitude = volume frequency = pitch

Vocaloid:

not sure about the equal signs, maybe there's another synonym for related you could use

zarkam21:

For a high loud note wave, there is a high freiquency as well as a high amplitude. The characteristics that are related to the amplitude and frequency are: amplitude is linked to volume frequency is linked to pitch

Vocaloid:

good

zarkam21:

Write two or three sentences to define intensity and to state how intensity changes as the distance from the sound is increased.

Vocaloid:

intensity = energy of a sound wave over a defined area perpendicular to the direction of propagation intensity is proportional to 1/distance^2 (inverse square law)

zarkam21:

Intensity is a sound wave energy over an assigned region. This is perpendicular to the path of propgation. Intensity changes as the distance from the sound is increased because of the inverse square law. This law tells that intensity is correspondent to 1/distance^2

Vocaloid:

good, just some minor tweaks * it's propagation not propgation * should say "intensity decreases" instead of just "intensity changes" to be more specific

zarkam21:

Calculate the energy of light that has a wavelength of 5.43 × 10–^7 m

Vocaloid:

E = hc/lambda where h = planck's constant, c = speed of light, lamda = wavelength

zarkam21:

0.543

zarkam21:

not sure if I did this right

Vocaloid:

I think that's a little off planck's constant = (6.62607×10^-34) speed of light = 3.0*10^8 lambda = (5.43 × 10^-7) therefore E = (6.62607×10^-34)*(3*10^8)/(5.43 × 10^-7) = ?

zarkam21:

3.66*10^-19

Vocaloid:

good, since I used the joules version of planck's constant the final answer will be in joules

zarkam21:

1 attachment
zarkam21:

I know that every color is absorbed but that color. But I am not explicit on this

Vocaloid:

well, one color is ~absorbed~ and one color is ~reflected~ (I think those are the "wave interactions" they're looking for?) which one is which?

zarkam21:

In this picture of wave interactions, since the blue light bounces off, it is reflected. However, since the red light does not, it is absorbed

Vocaloid:

A+ so to answer the second question, what color is the object? red or blue?

zarkam21:

The color of the object would be blue because it is the one that is reflected

Vocaloid:

A+ again that's at least two sentences so I believe that is sufficiently answered

zarkam21:

X-rays and visible light are parts of the electromagnetic spectrum. Write three or four sentences to compare these electromagnetic waves with respect to the following properties: wavelength, energy, frequency, and speed.

Vocaloid:

 wavelength, energy, and frequency are pretty straightforward by looking at an EM spectrum the only tricky one is speed, all electromagnetic radiation travels at the speed of light in a vacuum, so x-rays and visible light would travel at the same speed (given that the medium is the same)

Vocaloid:

actually wait let me modify that last part a bit

Vocaloid:

all electromagnetic radiation travels at the speed of light in a vacuum, so x-rays and visible light would travel at the same speed in a vacuum (i'm actually not sure about other media)

zarkam21:

To describe wavelength, the shorter the length the more compressed the wave. The longer the length. For energy, the higher the energy, the more compressed the wave, For frequency, the higher the frequency, the more compressed the wave. To compare x-rays and visible light, they travel at the speed of light a vacuum has

Vocaloid:

let's try from the beginning ^^ first sentence: compare the wavelengths of x-rays and gamma rays, which one is higher? second sentence: compare the energies, which one has a higher energy? third sentence: compare the frequencies, which one has higher frequency? fourth sentence: compare the speeds

zarkam21:

high energy?

zarkam21:

for the first one

Vocaloid:

first sentence: do x-rays or gamma rays have the higher wavelength?

zarkam21:

To compare wavelength between x-rays and gamma rays, x-rays have the higher wavelength. To compare energies between x-rays and gamma rays, gamma rays have the higher energy.To compare frequencies between x-rays and gamma rays, gamma rays have the higher frequency. To compare speed between x-rays and gamma rays, gamma rays have the higher speed.

Vocaloid:

good but check that last sentence again, gamma rays and x-rays have the same speed in a vacuum anyway I have to go, good luck with your studies, just be careful when you submit your answers b/c some of your posts on here might show up in the plagiarism checker

zarkam21:

Thank you so much!